OpenStudy (anonymous):

integral (sq rt of x to the 4th root -6x^2+e^x-2/x) dx

OpenStudy (loser66):

can you use draw box below to draw it out? I don't understand your problem

OpenStudy (anonymous):

\[\int\limits \left( \sqrt[4]{x}-6x^2 +e^x-2/x \right)\]

OpenStudy (anonymous):

I was looking for that to type my equation when i posted my question originally and couldnt find it

OpenStudy (loser66):

\[\int\limits \left( x^{\frac{1}{4}}-6x^2 +e^x-2x^{-1}\right)dx\] yeah, easy now, right?

OpenStudy (anonymous):

thats about as far as i got on the problem

OpenStudy (anonymous):

wasn't sure if i take the antiderivative next or not

OpenStudy (loser66):

\[=\frac{x^{\frac{1}{4}+1}}{\frac{1}{4}+1}-4x^3+e^x-2lnx+C\]

OpenStudy (loser66):

simplify the first term \[\frac{4}{5}*x^{\frac{5}{4}}-4x^3+e^x-2lnx+C\]

OpenStudy (loser66):

that's it

OpenStudy (anonymous):

ok, so i see you combined the two x's to get a 4 instead of a 12 like i did

OpenStudy (loser66):

I have class now, cannot explain . sorry for that.

OpenStudy (loser66):

@UnkleRhaukus please take it over,

OpenStudy (loser66):

@ganeshie8 Please

OpenStudy (anonymous):

i donno why but i got \[\frac{ 1 }{ 4 }x^{-3/4}-12x+e^x+2x ^{-2}\]

OpenStudy (unklerhaukus):

\[\int\limits \left( \sqrt[4]{x}-6x^2 +e^x-\frac2x \right)\,\mathrm dx\\ =\int\limits \left( x^{\frac{1}{4}}-6x^2 +e^x-2x^{-1}\right)\,\mathrm dx\\ =\int\limits x^{\frac{1}{4}}\,\mathrm dx-\int\limits6x^2\,\mathrm dx+\int\limits e^x\,\mathrm dx-2\int\frac {\mathrm dx}x\\ =\frac{x^{\frac{1}{4}+1}}{\frac{1}{4}+1}-\frac{6x^{2+1}}{2+1}+e^x-2\ln x+C\\=\]

OpenStudy (unklerhaukus):

you are differentiation where you should be integrating

OpenStudy (anonymous):

OH I see! Thank you

OpenStudy (unklerhaukus):

\[\frac{\mathrm d}{\mathrm dx}(x^n+C)=nx^{n-1}\\----------\\ \int x^n\mathrm dx=\frac{x^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

ok thank you!

OpenStudy (unklerhaukus):

you notice that if the index \(n\) is -1, \[\begin{align}\int x^{-1}\mathrm dx&=\color{red}{\cancel{\color{black}{\frac{x^{-1+1}}{-1+1}+C =\frac{1}{0}+C???}}}\\ \\&=\ln x+C\\ \end{align}\] it dosent work, you have to

OpenStudy (unklerhaukus):

... know this .